OpenStudy (calculusxy):

How many cubes of edge length 1/4 inch are required to completely pack the rectangular prism without any gaps or overlap? (I will draw the cube)

OpenStudy (calculusxy):



OpenStudy (calculusxy):

@wio

OpenStudy (anonymous):

Okay, so let me explain more generally. \[ \left(\frac{1}{4}\text{in}\right)^3 = \frac{1}{64}\text{in}^3 \]

OpenStudy (anonymous):

So all you have to do is find the volume and multiply by \(64\).

OpenStudy (calculusxy):

Okay so when I found the volume, I got 27/8. So I multiply 27/8*64/1?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

Thanks. Can you help me with more?